DIFFERENTIAL GEOMETRY/PDE SEMINAR

Friday, May 19 2:45 PM in SMITH 107 (Note: special day, time and room)

Strong unique continuation for second order elliptic equations

Daniel Tataru (Northwestern University)

We present recent sharp unique continuation results and counterexamples for second order elliptic equations with nonsmooth coefficients and unbounded potentials.